# SuperOcc: Toward Cohesive Temporal Modeling for Superquadric-based Occupancy Prediction

### v1

@@ I Introduction
3D occupation prediction is essentially a 3D semantic segmentation task. Therefore, early methods [21, 43, 48, 39, 38] directly generate dense volumetric feature representations followed by per-voxel classification. These methods are intuitive and capable of capturing sufficient geometric details, but they suffer from high computational complexity and memory consumption. To alleviate this burden, subsequent studies have extensively explored alternative dense representations [45, 46, 13] and more efficient learning schemes [42, 32]. Although these methods effectively reduce computational costs, they still fail to fully exploit the spatial sparsity of real-world scenes, i.e., the vast majority of 3D space is actually free. Consequently, performing uniform computation and prediction across these free regions introduces considerable computational redundancy and resource waste.
@@ II-A1 Dense Scene Representation
Dense scene representation is the most straightforward modeling paradigm. Existing methods [2, 21, 43, 48, 38, 32, 42, 19, 34, 16] typically discretizes the 3D space into voxel grids to construct a dense volume representation. FB-OCC [21] employs depth-aware back-projection to complete and enhance the initial 3D volume generated by forward projection. SurroundOcc [43] adopts a 2D-3D spatial attention mechanism to integrate multi-camera information into 3D volume queries and construct 3D volume features in a multi-scale manner. Although these methods are capable of capturing rich geometric details, they suffer from high computational cost and memory overhead. Subsequent works therefore aim to improve the efficiency of dense volume construction and processing. OccFormer[48] decomposes heavy 3D processing into local and global transformer pathways along the horizontal plane to efficiently process the 3D volume. CTF-Occ [38] introduces an incremental token selection strategy to mitigate the computational burden caused by dense sampling. To better balance performance and efficiency, COTR [32], PanoOcc [42] first generate a compact low-resolution volume and then recover the geometric details via voxel upsampling.
@@ II-B Camera-based Temporal Modeling
The view-centric paradigm extracts historical information from structured scene representations across frames, such as perspective view (PV), bird’s-eye-view (BEV), or 3D volume representations. Specifically, PV-based methods [30, 28, 24, 27, 40] typically perform temporal modeling by interacting multi-frame PV features with object queries. For example, PETRv2 [30] utilizes 3D position embedding to transform multi-frame, multi-view 2D features into 3D position-aware features, and then object queries interact with these features using global cross-attention. To improve the interaction efficiency, SparseBEV [28], Sparse4D [24], SparseOcc [27], OPUS [40] adopt adaptive sparse sampling and aggregation. In parallel, BEV/volume-based methods [20, 36, 42, 9, 4, 5, 18, 17, 49, 44, 3, 6] focus on the fusion of multi-frame BEV/volume features. SOLOFusion [36], PanoOcc [42], GSD-Occ [9] and ALOcc [4] align historical features to the current frame based on ego-motion, and then fuse them with current features via feature concatenation. BEVFormer [20] and ViewFormer [18] utilize temporal self-attention to extract temporal information from history BEV features.
@@ III-A Preliminaries: 3D Semantic Superquadric Representation
$\mathcal{S}=\{\mathbf{S}^{i}\}_{i=1}^{M}=\left\{\bigl(\textbf{m}^{i},\textbf{r}^{i},\textbf{s}^{i},\mathbf{\epsilon}^{i},\sigma^{i},\textbf{c}^{i}\bigr)\right\}_{i=1}^{M},$ (1)
where the center $\mathbf{m}=(m_{x},m_{y},m_{z})^{T}$ and rotation $\mathbf{r}\in\mathbb{R}^{4}$ specify the pose of the superellipsoid. $\mathbf{s}=(s_{x},s_{y},s_{z})^{T}$ determines the scaling factors along the three axes, while the squareness parameter $\mathbf{\epsilon}=(\epsilon_{1},\epsilon_{2})^{T}$ controls the squareness of the shape. Opacity $\sigma$ and semantic probability $\mathbf{c}$ are equipped to incorporate semantic information.
@@
$p_{s}(\mathbf{p})=\frac{\sum_{i=1}^{M}p_{o}(\mathbf{p};\mathbf{S}^{i})\,\sigma^{i}\mathbf{c}^{i}}{\sum_{j=1}^{M}p_{o}(\mathbf{p};\mathbf{S}^{j})\sigma^{j}}.$ (5)
@@
$o(\mathbf{p})=[p_{o}(\mathbf{p})\cdot p_{s}(\mathbf{p});1-p_{o}(\mathbf{p})],$ (6)
@@ III-B Overall framework
As illustrated in Fig. 2, SuperOcc follows an encoder-decoder architecture. The image encoder comprises a backbone (e.g., ResNet50 [8]) and a neck (e.g., FPN [23]). Given multi-view images at time $t$ as input, the image encoder extracts multi-scale features denoted as $I^{t}=\{I^{t,v,l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}|1\leq l\leq L,1\leq v\leq V\}$, where $V$ and $L$ represent the number of camera views and feature scales, respectively. $H_{l}$ and $W_{l}$ denote the feature resolution at the $l$-th feature scale. To leverage temporal information, the image features from the most recent $T$ frames are maintained in a memory queue, denoted as $\mathcal{I}=\{I^{t}\}_{t=t_{0}-T}^{t_{0}}$.
@@ III-C1 View-centric Temporal Modeling
Specifically, for each query $\mathbf{q}$ and its corresponding 3D reference point $\mathbf{p}_{ref}=(x,y,z)$, we first utilize a linear layer to predict a set of sampling offsets:
@@
$\mathcal{P}_{s}=\{\mathbf{p}^{i}=(x+\triangle x^{i},y+\triangle y^{i},z+\triangle z^{i})\}_{i=1}^{N_{s}}.$ (8)
@@
To sample features from any time $t$ within the memory queue $\mathcal{I}$, we project each sampling point onto the hit camera views based on ego-motion compensation and camera parameters. The sampled feature $f^{t,i}$ for the $i$-th point at time $t$ is extracted via bilinear interpolation as follows:
$f^{t,i}=\frac{1}{|V_{hit}|}\sum\limits_{v\in V_{hit}}\sum\limits_{l=1}^{L}w^{i,l}\cdot\mathcal{BS}(I^{t,v,l},\pi_{v}(T_{t_{0}\rightarrow{t}}p^{i})),$ (9)
where $V_{hit}$ is the set of hit camera views, $T_{t_{0}\rightarrow{t}}\in\mathbb{R}^{4\times 4}$ denotes the ego-motion transformation matrix that transforms the 3D coordinates from the current coordinate frame at time $t_{0}$ to the target frame at time $t$. $\pi_{v}(\cdot)$ denotes the projection function onto the $v$-th image plane, and $\mathcal{BS}(\cdot)$ denotes the bilinear sampling function. The term $w^{i,l}$ represents the weight predicted by the query for the $i$-th sampling point on the $l$-th feature scale.
@@ III-C2 Object-centric Temporal Modeling
At each frame, the top-$N_{p}$ queries with the highest foreground scores and their reference points $P_{p}$ are pushed into a memory queue. After temporal alignment, they are propagated to the next frame. To alleviate potential spatial redundancy between propagated queries and newly initialized ones, we discard initialized queries whose reference points lie within a distance threshold $\tau$ (e.g., $\tau=1.0\text{ m}$) of any propagated reference point. Then, the query set is complemented by stochastically sampling $N_{q}-N_{p}$ candidates from the residual initialization queries to maintain a constant budget $N_{q}$. This ensures that the hybrid queries provide both stable temporal priors for persistent objects and sufficient exploratory coverage for newborn objects.
@@ III-D Multi-Superquadric Decoding Strategy
$\mathcal{S}_{local}=\{\mathbf{S}^{k}\}_{k=1}^{K}=\{(\mathbf{p}_{ref}+\mathbf{o}^{k},\mathbf{r}^{k},\mathbf{s}^{k},\mathbf{\epsilon}^{k},\sigma^{k},\mathbf{c})\bigr\}_{k=1}^{K},$
@@
$\{(\mathbf{o}^{k},\mathbf{r}^{k},\mathbf{s}^{k},\mathbf{\epsilon}^{k},a^{k})\}_{k=1}^{K}=\phi_{\text{reg}}(\mathbf{q}).$ (12)

### v2

# SuperOcc: Toward Cohesive Temporal Modeling for Superquadric-based 3D Occupancy Prediction

@@ I Introduction
3D occupancy prediction is essentially a 3D semantic segmentation task. Therefore, early methods [21, 45, 50, 41, 40] directly generate dense volumetric feature representations followed by per-voxel classification. These methods are intuitive and capable of capturing sufficient geometric details, but they suffer from high computational complexity and memory consumption. To alleviate this burden, subsequent studies have extensively explored alternative dense representations [47, 48, 13] and more efficient learning schemes [44, 32]. Although these methods effectively reduce computational costs, they still fail to fully exploit the spatial sparsity of real-world scenes, i.e., the vast majority of 3D space is actually free. Consequently, performing uniform computation and prediction across these free regions introduces considerable computational redundancy and resource waste.
@@ II-A1 Dense Scene Representation
Dense scene representation is the most straightforward modeling paradigm. Existing methods [2, 21, 45, 50, 40, 32, 44, 19, 34, 16, 39, 37] typically discretizes the 3D space into voxel grids to construct a dense volume representation. FB-OCC [21] employs depth-aware back-projection to complete and enhance the initial 3D volume generated by forward projection. SurroundOcc [45] adopts a 2D-3D spatial attention mechanism to integrate multi-camera information into 3D volume queries and construct 3D volume features in a multi-scale manner. Although these methods are capable of capturing rich geometric details, they suffer from high computational cost and memory overhead. Subsequent works therefore aim to improve the efficiency of dense volume construction and processing. OccFormer[50] decomposes heavy 3D processing into local and global transformer pathways along the horizontal plane to efficiently process the 3D volume. CTF-Occ [40] introduces an incremental token selection strategy to mitigate the computational burden caused by dense sampling. To better balance performance and efficiency, COTR [32], PanoOcc [44] first generate a compact low-resolution volume and then recover the geometric details via voxel upsampling.
@@ II-B Camera-based Temporal Modeling
The view-centric paradigm extracts historical information from structured scene representations across frames, such as perspective view (PV), bird’s-eye-view (BEV), or 3D volume representations. Specifically, PV-based methods [30, 28, 24, 27, 42] typically perform temporal modeling by interacting multi-frame PV features with object queries. For example, PETRv2 [30] utilizes 3D position embedding to transform multi-frame, multi-view 2D features into 3D position-aware features, and then object queries interact with these features using global cross-attention. To improve the interaction efficiency, SparseBEV [28], Sparse4D [24], SparseOcc [27], OPUS [42] adopt adaptive sparse sampling and aggregation. In parallel, BEV/volume-based methods [20, 36, 44, 9, 4, 5, 18, 17, 51, 46, 3, 6] focus on the fusion of multi-frame BEV/volume features. SOLOFusion [36], PanoOcc [44], GSD-Occ [9] and ALOcc [4] align historical features to the current frame based on ego-motion, and then fuse them with current features via feature concatenation. BEVFormer [20] and ViewFormer [18] utilize temporal self-attention to extract temporal information from historical BEV features.
@@
However, these two paradigms are typically explored independently. View-centric methods capture fine-grained spatio-temporal context, whereas object-centric methods efficiently propagate historical priors. To leverage their complementary strengths, we propose a cohesive temporal modeling mechanism that jointly exploits view-centric and object-centric temporal cues within a unified framework.
@@ III-A Preliminaries: 3D Semantic Superquadric Representation
$\mathcal{S}=\{\mathbf{S}^{i}\}_{i=1}^{M}=\left\{\bigl(\textbf{m}^{i},\textbf{r}^{i},\textbf{s}^{i},\bm{\epsilon}^{i},\sigma^{i},\textbf{c}^{i}\bigr)\right\}_{i=1}^{M},$ (1)
where the center $\mathbf{m}=(m_{x},m_{y},m_{z})^{T}$ and rotation $\mathbf{r}\in\mathbb{R}^{4}$ specify the pose of the superellipsoid. $\mathbf{s}=(s_{x},s_{y},s_{z})^{T}$ determines the scaling factors along the three axes, while the squareness parameter $\bm{\epsilon}=(\epsilon_{1},\epsilon_{2})^{T}$ controls the squareness of the shape. Opacity $\sigma$ and semantic probability $\mathbf{c}$ are equipped to incorporate semantic information.
@@
$\mathbf{p_{s}}(\mathbf{p})=\frac{\sum_{i=1}^{M}p_{o}(\mathbf{p};\mathbf{S}^{i})\,\sigma^{i}\mathbf{c}^{i}}{\sum_{j=1}^{M}p_{o}(\mathbf{p};\mathbf{S}^{j})\sigma^{j}}.$ (5)
@@
$\mathbf{o}(\mathbf{p})=[p_{o}(\mathbf{p})\cdot\mathbf{p_{s}}(\mathbf{p});1-p_{o}(\mathbf{p})],$ (6)
@@ III-B Overall framework
As illustrated in Fig. 2, SuperOcc follows an encoder-decoder architecture. The image encoder comprises a backbone (e.g., ResNet50 [8]) and a neck (e.g., FPN [23]). Given multi-view images at time $t$ as input, the image encoder extracts multi-scale features denoted as $I^{t}=\{I^{t,v,l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}|1\leq l\leq L,1\leq v\leq V\}$, where $V$, $L$, and $C$ represent the number of camera views, feature scales, and feature channels, respectively. $H_{l}$ and $W_{l}$ denote the feature resolution at the $l$-th feature scale. To leverage temporal information, the image features from the most recent $T$ frames are maintained in a memory queue, denoted as $\mathcal{I}=\{I_{t},I_{t-1},\cdots,I_{t-T+1}\}$.
@@ III-C1 View-centric Temporal Modeling
Specifically, for each query $\mathbf{q}$ and its corresponding 3D reference point $\mathbf{p}_{ref}=(x,y,z)^{T}$, we first utilize a linear layer to predict a set of sampling offsets:
@@
$\mathcal{P}_{s}=\{\mathbf{p}^{i}=(x+\triangle x^{i},y+\triangle y^{i},z+\triangle z^{i})^{T}\}_{i=1}^{N_{s}}.$ (8)
@@
To sample features from any time $\tau\in\{t,t-1,\cdots,t-T+1\}$ within the memory queue $\mathcal{I}$, we project each sampling point onto the hit camera views based on ego-motion compensation and camera parameters. The sampled feature $f^{\tau,i}$ for the $i$-th point at time $\tau$ is extracted via bilinear interpolation as follows:
$f^{\tau,i}=\frac{1}{|V_{hit}|}\sum\limits_{v\in V_{hit}}\sum\limits_{l=1}^{L}w^{i,l}\cdot\mathcal{BS}(I^{\tau,v,l},\pi_{v}(T_{t\rightarrow{\tau}}\mathbf{p}^{i})),$ (9)
where $V_{hit}$ is the set of hit camera views, $T_{t\rightarrow{\tau}}\in\mathbb{R}^{4\times 4}$ denotes the ego-motion transformation matrix that transforms the 3D coordinates from the coordinate frame at current time $t$ to the target frame at time $\tau$. $\pi_{v}(\cdot)$ denotes the projection function onto the $v$-th image plane, and $\mathcal{BS}(\cdot)$ denotes the bilinear sampling function. The term $w^{i,l}$ represents the weight predicted by the query for the $i$-th sampling point on the $l$-th feature scale.
@@ III-C2 Object-centric Temporal Modeling
At each frame, the top-$N_{p}$ queries with the highest foreground scores and their reference points $P_{p}$ are pushed into a memory queue. After temporal alignment, they are propagated to the next frame. To alleviate potential spatial redundancy between propagated queries and newly initialized ones, we discard initialized queries whose reference points lie within a distance threshold $\delta$ (e.g., $\delta=1.0\text{ m}$) of any propagated reference point. Then, the query set is complemented by stochastically sampling $N_{q}-N_{p}$ candidates from the residual initialization queries to maintain a constant budget $N_{q}$. This ensures that the hybrid queries provide both stable temporal priors for persistent objects and sufficient exploratory coverage for newborn objects.
@@ III-D Multi-Superquadric Decoding Strategy
$\mathcal{S}_{local}=\{\mathbf{S}^{k}\}_{k=1}^{K}=\{(\mathbf{p}_{ref}+\mathbf{o}^{k},\mathbf{r}^{k},\mathbf{s}^{k},\bm{\epsilon}^{k},\sigma^{k},\mathbf{c})\bigr\}_{k=1}^{K},$
@@
$\{(\mathbf{o}^{k},\mathbf{r}^{k},\mathbf{s}^{k},\bm{\epsilon}^{k},a^{k})\}_{k=1}^{K}=\phi_{\text{reg}}(\mathbf{q}).$ (12)